# Fluctuations of the free energy of the spherical Sherrington–Kirkpatrick model with heavy-tailed interaction

### v1

@@ Abstract
We consider the 2-spin spherical Sherrington–Kirkpatrick model where the interactions between the spins are given as random variables with heavy-tailed distribution. We prove that the free energy exhibits sharp phase transition, depending on the location of the largest eigenvalue of the interaction matrix. We also prove the order of the limiting free energy and the limiting distribution of the fluctuation of the free energy for both regimes.
@@ 1 Introduction
A notable variant of the SK model is the spherical SK model, introduced by Kosterlitz, Thouless, and Jones [17]. It replaces the standard spin vectors with those uniformly distributed on a sphere. The spherical SK model has garnered attention for its applications in statistical mechanics and other fields like information theory [21], with significant progress made in understanding its mathematical properties, including its limiting free energy [25] and the fluctuation of the free energy [5]. For other variants of the SSK model, we refer to [4] and references therein.
While the exact formulas for the limit and the fluctuation of the free energy may depend on the detail of the model, we can find one common feature in these models: The limiting free energy is a continuous function of the inverse temperature. Technically, it can be understood as follows: (1) by the integral representation formula (e.g., Lemma 1.3 in [5]), the partition function depends only on the eigenvalues of the interaction matrix, (2) the eigenvalues are close to the deterministic locations determined by their limiting empirical spectral distribution (ESD), (3) as a result, the limiting free energy can be well-approximated by a deterministic function of the inverse temperature.
In this paper, we consider a variant of the SSK model where the interactions between the spins are given as random variables with heavy-tailed distribution and investigate how this heavy-tailed interaction can alter the continuity of the limiting free energy. Indeed, there are several reasons that we can believe that the heavy-tailed SSK model can exhibit a feature that is significantly different from other SSK models. First, the ESDs of the heavy-tailed random matrices are markedly different from corresponding ‘light-tailed’ Wigner random matrices. Second, recent advancements have been made in calculating the free energy fluctuation for the SK model with heavy-tailed interactions [10], revealing significant differences from the usual SK model.
Before describing our model more in detail, we recall the key results for the usual SSK model [5]. The Hamiltonian for the SSK model is defined by
@@
for a real symmetric random matrix $H$, where $\sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{N})\in S_{N}:=\{x\in\mathbb{R}^{N}:x_{1}^{2}+\dots+x_{N}^{2}=N\}$. The entries of $H$ are with mean 00 and variance $1$, so that the largest eigenvalue of $H$ is of order $\sqrt{N}$ with high probability. The partition function and the free energy of the SSK model is defined as
@@
In this paper, we further assume that the entries $H$ are heavy-tailed random variables in the sense that $\mathbb{P}(|H_{ij}|>u)\sim u^{-\alpha}$ for some $\alpha\in(0,2)$. The exponent $\alpha$ determines the heaviness of the tail of the distributions. With heavy-tailed entries, unlike the light-tailed model, the largest eigenvalue is not of order $1$ but grows with $N$. We thus need to introduce an additional normalization factor $b_{N}$, which is order asymptotically $N^{2/\alpha}$, and consider the partition function defined as
@@
See Section 2.1, especially Definition 2.5, for the precise definition of the model and the partition function $Z_{N}$ with the normalization factor $b_{N}$. We remark that the normalization factor $b_{N}$ ensures that the extreme eigenvalues of $b_{N}^{-1}H$ is of order $1$; see [3, 24] for more detail.
@@ 1.1 Main contribution
When compared to the behavior of the free energy for the usual SSK model, we notice several contrasts. First, while for the usual SSK model the size of $\log Z_{N}$ is of order $N$ regardless of the temperature, it is markedly different for our model in the high temperature regime and the low temperature regime - the former is of order $1$ whereas the latter is of order $N$. Second, for the usual SSK model the fluctuation of $\log Z_{N}$ is much smaller than its deterministic limit, they are of comparable order for our model.
The transition between the high temperature regime and the low temperature regime is even more striking for our model. For the usual SSK model, the behavior of the free energy is solely determined by the inverse temperature. It is in the high temperature regime if $\beta$ is below a certain threshold, $1/2$ in (1.3), and in the low temperature regime if $\beta$ is below the threshold; no randomness is involved in the phase transition. On the other hand, with heavy-tailed interaction matrix, the distinction between the two phases are more subtle. In Theorem 2.6, we prove that the high temperature regime is when the largest eigenvalue of the interaction matrix is above a certain threshold determined by $\beta$ and the low temperature regime is when the largest eigenvalue is below the threshold. It in particular implies that the size of $\log Z_{N}$ is of order $1$ with probability $P_{\beta}$ and of order $N$ with probability $1-P_{\beta}$. The change of the inverse temperature only changes the probability $P_{\beta}$.
@@
The main technical difficulty in the analysis is due to that some spectral properties of heavy-tailed matrices are not known, e.g., its LSS with respect to a logarithmic function. To handle it, we notice that the fluctuation of several quantities such as the trace and the Hilbert-Schmidt norm of a heavy-tailed matrix is governed by a few extreme entries of the matrix. Then, by applying well-known facts about heavy-tailed distributions, we can analyze the LSS with respect to a logarithmic function that is essential in the analysis for high temperature regime.
@@ 1.2 Related works
The SSK model was first introduced by Kosterlitz, Thouless, and Jones [17]. The limiting free energy for the SSK model was computed (non-rigorously) in [17] and proved by Talagrand [25]. The fluctuation of the free energy for the SSK model was proved in [5]. Corresponding results for several variants of the SSK model were proved in [5, 6, 7, 20, 16]. Several other results were also proved, including the near-critical behavior [8, 18, 15] and the overlap [19]. For more results on the SSK model and its variants, we refer to [4] and references therein.
The spectral properties of the heavy-tailed random matrices have been extensively studied in random matrix theory. The behavior of the largest eigenvalue was first proved by Soshnikov [24] for the case $\alpha<2$ and later generalized to the case $\alpha<4$ by Auffinger, Ben Arous, and Péché [3]. Several other results were proved, including the spectrum normalization [2], a central limit theorem for the Stieltjes transform [9], and bulk universality [1].
@@ 2.1 Definitions
We say that a random variable $X$ is heavy-tailed with exponent $\alpha$ if $\mathbb{P}(|X|>u)=L(u){u^{-\alpha}}$ for some $\alpha\in(0,1)$ and some slowly varying function $L$, i.e. $\lim_{x\to\infty}L(tx)/L(x)=1$ for all $t$. We further assume that for any $\delta>0$, there exists $x_{0}$ such that $L(x)e^{x^{\delta}}$ is an increasing function on $(x_{0},\infty)$.
@@
Our model is spherical SK model in which the interactions between the spins are i.i.d. heavy-tailed random variables. Its partition function and the free energy are defined as follows:
@@
$b_{N}=\inf\{t:\mathbb{P}(|M_{11}|>t)\leq\frac{2}{N(N+1)}\}$ (2.1)
where $L$ is the slowly varying function associated the entries of $M$ as in Definition 2.2. We define the partition function associated with $M$ at inverse temperature $\beta$ by
@@ 2.2 Main results
$\mathbb{P}(X>u)=\exp(-x^{-\alpha}+(2\beta)^{\alpha}).$
@@
See Proposition 4.1 in Section 4 and Proposition 5.1 in Section 5. ∎
@@ 2.3 Strategy of the proof
For the proof of Theorem 2.6, we adapt the integral representation formula for the partition function, introduced in [17] and proved in [5]. Let $\lambda_{1}\geq\lambda_{2}\geq\dots\lambda_{N}$ be the eigenvalues of $M$. In Lemma 3.16, we prove that
@@
While $G^{\prime}(z)$ is a random function, since $\lambda_{i}=O(1)$ for most of the eigenvalues $\lambda_{i}$, it is possible to approximate $G^{\prime}$ by a deterministic function with negligible error. We can then approximate the solution of the equation (2.6) by a deterministic number $b_{N}/(2\beta)$. Thus, if $\lambda_{1}<b_{N}/(2\beta)$, which we call the high temperature regime, we can analyze the asymptotic behavior of $Z_{N}$ by choosing $\gamma=b_{N}/(2\beta)$. In Section 4, we consider this case and prove the result on the fluctuation of $\log Z_{N}$, Proposition 4.1, by considering the linear spectral statistics of $M$ involving the function $G(z)$ in (2.5).
@@ 3.1 Properties of heavy-tailed Wigner matrices
In this section, we collect several important results on heavy-tailed random variables and heavy-tailed Wigner matrices. For the sake of convenience, throughout this section we assume that $M$ is an heavy-tailed Wigner matrix with exponent $\alpha$ as in Definition 2.4 and each entry $M_{ij}$ satisfies $\mathbb{P}(|M_{ij}|>u)=G(u)=L(u)u^{-\alpha}$, where $L$ is a slowly varying function as in Definition 2.2. We let $\lambda_{1}\geq\dots\geq\lambda_{N}$ be the eigenvalues of $M$. We first define the high probability events.
@@
We say that an event $F_{N}$ holds with high probability if there exist (small) $\delta>0$ and (large) $N_{0}$ such that $\mathbb{P}(F_{N})<N^{-\delta}$ holds for any $N>N_{0}$.
@@
For the proof of Proposition 3.2, see Theorem 1 in [24]. Next, we state the results that ensure good events for $M$ hold with high probability.
@@
For every $\epsilon>0$, with high probability, that the following holds for any $i=1,2,\dots,N$: Suppose that $|M_{ii^{\prime}}|=\max_{j}|M_{ij}|$. Then, $|M_{ii^{\prime}}|<b_{N}^{1/2+\epsilon}$ or $|\sum_{j:j\neq i^{\prime}}M_{ij}|<b_{N}^{1/2+\epsilon}$.
@@
For every $\epsilon>0$, these three conditions hold with high probability.
  • $\lambda_{1}>b_{N}N^{-\epsilon}$
  • $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$
  • If $\lambda_{1}<Cb_{N}$, then $\lambda_{1}<Cb_{N}-b_{N}N^{-\epsilon}$
@@
In Proposition 3.7, we can actually prove that both $c_{N}/a_{N}$ and $S_{N}/a_{N}$ converge. To see this, for $\alpha<1$, we consider
@@
where $g$ is the probability density function for $|X_{1}|$ and $G(x)=\mathbb{P}(|X_{1}|>x)$. The last term in the right side of the equation above is then
$\int_{a}^{a_{N}}G(x)\,\mathrm{d}x=\int x^{-\alpha}L(x)\,\mathrm{d}x=\frac{1}{1-\alpha}a_{N}^{-\alpha+1}L(a_{N})(1+o(1))=\frac{1}{1-\alpha}\frac{a_{N}}{N}(1+o(1))$
by Karamata’s theorem. (See, e.g., Thm 8.9.2 of [13]). Thus
$N\mathbb{E}[|X_{1}|\mathbbm{1}_{|X_{1}|\leq a_{N}}]/a_{N}\to\frac{1}{1-\alpha}-1,$
which implies that $c_{N}/a_{N}$ converges in probability to a constant depending only on the probability density function. Together with Proposition 3.7, this also shows that $S_{N}/a_{N}$ converges.
@@
See Appendix A for the proof. Lemma below is just corollary of this lemma.
@@
See Appendix A for the proof. The next lemma provides the information about the asymptotic for the number of eigenvalues that is comparable with the largest eigenvalue.
@@
The eigenvlaues $\lambda_{1},..,\lambda_{N}$ of M satisfy
@@
Furthermore, we also need the limit of $\sum_{i}\lambda_{i}/b_{N}$ and it converges to 0 as below lemma.
@@
The eigenvlaues $\lambda_{1},..,\lambda_{N}$ of $M$ satisfies

### v2

@@ Abstract
We consider the 2-spin spherical Sherrington–Kirkpatrick model without external magnetic field where the interactions between the spins are given as random variables with heavy-tailed distribution. We show that the free energy exhibits a sharp phase transition depending on the location of the largest eigenvalue of the interaction matrix. We also prove the order of the limiting free energy and the limiting distribution of the fluctuation of the free energy for both regimes.
keywords:Heavy-tailed matrix, SSK model, Free energy, Phase transition
@@ 1 Introduction
A notable variant of the SK model is the spherical SK model, introduced by Kosterlitz, Thouless, and Jones [25]. Their model replaces the standard spin vectors with those uniformly distributed on a sphere. The spherical SK model has garnered attention for its applications in statistical mechanics and other fields such as information theory [29], with significant progress made in understanding its mathematical properties, including its limiting free energy [34] and the fluctuation of the free energy [5]. For other variants of the SSK model, we refer to [4] and references therein.
Although the exact formulas for the limit and the fluctuation of the free energy may depend on the details of the model, we can find one common feature in these models: The limiting free energy is a deterministic function of the inverse temperature. Technically, it can be understood as follows: (1) by the integral representation formula (e.g., Lemma 1.3 in [5]), the partition function depends only on the eigenvalues of the interaction matrix, (2) the eigenvalues are close to the deterministic locations determined by their limiting empirical spectral distribution (ESD), (3) as a result, the limiting free energy can be well-approximated by a deterministic function of the inverse temperature.
In this paper, we consider a variant of the SSK model without external magnetic field in which the interactions between the spins are given as random variables with heavy-tailed distribution and investigate how this heavy-tailed interaction can alter the continuity of the limiting free energy. The non-spherical counterpart of such a model, i.e., the SK model with heavy-tailed interaction, was considered in [13, 14] and its properties including the phase transitions of several thermodynamic quantities were analyzed by replica method in [22, 21, 30]. Recently, the existence of free energy and its fluctuation in the high-temperature regime for the heavy-tailed SK model was proved in [20, 11] and its disordered chaos phenomenon at zero temperature was shown in [12]. These results reveal the significant differences between the heavy-tailed SK model and the usual SK model. While there might be several reasons that the heavy-tailed SK or SSK model is different from other models, we can consider the following: (1) the ESDs of the heavy-tailed random matrices are markedly different from corresponding ‘light-tailed’ Wigner random matrices, and (2) the conventional techniques such as the integration by parts cannot be used due to the unboundedness of the first a few moments (most importantly, the variance), unlike the analysis for the usual SK model, e.g., the proof of the existence of the thermodynamic limit for the free energy in [19].
Before describing the precise detail of our model, we recall the key results for the usual SSK model [5]. The Hamiltonian for the SSK model is defined by
@@
for a real symmetric random matrix $H$, where $\sigma=(\sigma_{1},\sigma_{2},\dots,\sigma_{N})\in S_{N}:=\{x\in\mathbb{R}^{N}:x_{1}^{2}+\dots+x_{N}^{2}=N\}$. The entries of $H$ are with mean 00 and variance $1$, so the largest eigenvalue of $H$ is of order $\sqrt{N}$ with high probability. The partition function and its free energy of the SSK model is defined as
@@
In our model, we further assume that the entries $H$ are heavy-tailed random variables in the sense that $\mathbb{P}(|H_{ij}|>u)\sim u^{-\alpha}$ for some $\alpha\in(0,2)$. The exponent $\alpha$ determines the heaviness of the tail of the distributions. With heavy-tailed entries, unlike the light-tailed model, the largest eigenvalue is not of the order $1$ but grows with $N$. We thus need to introduce an additional normalization factor $b_{N}$, which is asymptotically of order $N^{2/\alpha}$, and consider the partition function defined as
@@
See Section 2.1, especially Definition 2.5, for the precise definition of the model and the partition function $Z_{N}$ with the normalization factor $b_{N}$. We note that the normalization factor $b_{N}$ ensures that the extreme eigenvalues of $b_{N}^{-1}H$ are of order $1$; see [3, 33] for more details.
@@ 1.1 Main contribution
When compared to the behavior of the free energy for the usual SSK model, we notice several contrasts. First, while for the usual SSK model the size of $\log Z_{N}$ is of order $N$ regardless of the temperature, it is markedly different for our model in the high temperature regime and the low temperature regime - the former is of order $1$ whereas the latter is of order $N$. Second, for the usual SSK model, the fluctuation of $\log Z_{N}$ is much smaller than its deterministic limit; they are of comparable order for our model.
The transition between the high temperature regime and the low temperature regime is even more striking for our model. For the usual SSK model, the behavior of the free energy is solely determined by the inverse temperature. It is in the high temperature regime if $\beta$ is below a certain threshold, $1/2$ in (1.3), and in the low temperature regime if $\beta$ is below the threshold; no randomness is involved in the phase transition. On the other hand, with heavy-tailed interaction matrix, the distinction between the two phases is more subtle. In Theorem 2.6, we prove that the high temperature regime is when the largest eigenvalue of the interaction matrix is above a certain threshold determined by $\beta$ and the low temperature regime is when the largest eigenvalue is below the threshold. In particular, it implies that the size of $\log Z_{N}$ is of order $1$ with probability $P_{\beta}$ and of order $N$ with probability $1-P_{\beta}$. The change in inverse temperature only changes probability $P_{\beta}$.
@@
The main technical difficulty in the analysis is due to that some spectral properties of heavy-tailed matrices are not known, e.g., its LSS with respect to a logarithmic function. To handle it, we notice that the fluctuation of several quantities, such as the trace and the Hilbert-Schmidt norm of a heavy-tailed matrix is governed by a few extreme entries of the matrix. Then, by applying well-known facts about heavy-tailed distributions, we can analyze the LSS with respect to a logarithmic function that is essential in the analysis for high temperature regime.
@@ 1.2 Related works
The SSK model was first introduced by Kosterlitz, Thouless, and Jones [25]. The limiting free energy for the SSK model was computed (nonrigorously) in [25] and proved by Talagrand [34]. The fluctuation of the free energy for the SSK model was proved in [5]. Corresponding results for several variants of the SSK model were proved in [5, 6, 7, 28, 24]. Several other results were also shown, including the near-critical behavior [8, 26, 23] and the overlap [27]. For more results on the SSK model and its variants, we refer to [4] and references therein.
The spectral properties of the heavy-tailed random matrix have been extensively studied in random matrix theory. The behavior of the largest eigenvalue was first proved by Soshnikov [33] for the case $\alpha<2$ and later generalized to the case $\alpha<4$ by Auffinger, Ben Arous, and Péché [3]. Several other results were proved, including the spectrum normalization [2], a central limit theorem for the Stieltjes transform [10], and bulk universality [1].
@@ 2.1 Definitions
We say that a random variable $X$ is heavy-tailed with exponent $\alpha$ if $\mathbb{P}(|X|>u)=L(u){u^{-\alpha}}$ for some $\alpha\in(0,2)$ and some slowly varying function $L$, i.e. $\lim_{x\to\infty}L(tx)/L(x)=1$ for all $t$. We further assume that for any $\delta>0$, there exists $x_{0}$ such that $L(x)e^{x^{\delta}}$ is an increasing function on $(x_{0},\infty)$.
@@
Our model is a spherical SK model where the interactions between spins are i.i.d. heavy-tailed random variables. The partition function and free energy are defined as follows:
@@
$b_{N}=\inf\{t:\mathbb{P}(|M_{11}|>t)\leq\frac{2}{N(N+1)}\}.$ (2.1)
We define the partition function associated with $M$ at inverse temperature $\beta$ by
@@ 2.2 Main results
$\mathbb{P}(X>u)=\frac{1-\exp(-u^{-\alpha})}{1-\exp(-(2\beta)^{\alpha})}.$
@@
See Proposition 4.1 in Section 4 and Proposition 5.1 in Section 5. (See also Proposition 3.2 for about the limits of $\mathbb{P}(F_{1})$ and $\mathbb{P}(F_{2})$.) ∎
@@ 2.3 Strategy of the proof
For the proof of Theorem 2.6, we adapt the integral representation formula for the partition function, introduced in [25] and proved in [5]. Let $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{N}$ be the eigenvalues of $M$. In Lemma 3.16, we prove that
@@
If $\lambda_{1}<b_{N}/2\beta$, which we call the high temperature regime, since $\lambda_{i}\ll\lambda_{1}$ holds for all but $o(N)$ eigenvalues $\lambda_{i}$, it is possible to approximate $G^{\prime}(\gamma)$ by $2\beta-b_{N}/\gamma$ with negligible error. We can then approximate the solution of equation (2.6) using a deterministic number $b_{N}/(2\beta)$. Thus, we can analyze the asymptotic behavior of $Z_{N}$ by choosing $\gamma=b_{N}/(2\beta)$. In Section 4, we consider this case and prove the result on the fluctuation of $\log Z_{N}$, Proposition 4.1, by considering the linear spectral statistics of $M$ involving the function $G(z)$ in (2.5).
@@ 3.1 Properties of heavy-tailed Wigner matrices
In this section, we collect several important results on heavy-tailed random variables and heavy-tailed Wigner matrices, which will be used in the analysis of our model. (For more results on the eigenvalues of heavy-tailed Wigner matrices, we refer to [33, 9].) For the sake of convenience, throughout this section we assume that $M$ is an heavy-tailed Wigner matrix with exponent $\alpha$ as in Definition 2.4 and each entry $M_{ij}$ satisfies $\mathbb{P}(|M_{ij}|>u)=G(u)=L(u)u^{-\alpha}$, where $L$ is a slowly varying function as in Definition 2.2. We let $\lambda_{1}\geq\dots\geq\lambda_{N}$ be the eigenvalues of $M$. We first define the high probability events.
@@
We say that an event $F_{N}$ holds with high probability if there exist (small) $\delta>0$ and (large) $N_{0}$ such that $\mathbb{P}(F_{N})\geq 1-N^{-\delta}$ holds for any $N>N_{0}$.
@@
For the proof of Proposition 3.2, see Theorem 1 in [33]. Next, we state the results which ensure that the good events for $M$ hold with high probability.
@@
For every $\epsilon>0$, with high probability the following holds for any $i=1,2,\dots,N$: Suppose that $|M_{ii^{\prime}}|=\max_{j}|M_{ij}|$. Then, $|M_{ii^{\prime}}|<b_{N}^{1/2+\epsilon}$ or $|\sum_{j:j\neq i^{\prime}}M_{ij}|<b_{N}^{1/2+\epsilon}$.
@@
For every $\epsilon>0$, the following hold with high probability.
  • $\lambda_{1}>b_{N}N^{-\epsilon}$.
  • $\lambda_{1}-\lambda_{2}>b_{N}N^{-\epsilon}$.
  • $\lambda_{1}\notin(\frac{b_{N}}{2\beta}(1-N^{-\epsilon}),\frac{b_{N}}{2\beta}(1+N^{-\epsilon}))$.
@@
In Proposition 3.7, we can actually prove that both $c_{N}/a_{N}$ and $S_{N}/a_{N}$ converge for $\alpha<1$. To see this, for $\alpha<1$, we consider
@@
where $g$ is the probability density function for $|X_{1}|$ and $G(x)=\mathbb{P}(|X_{1}|>x)$. From the definition of $a_{N}$, we find that $Na_{N}^{-\alpha}L(a_{N})=1+o(1)$. The last term in the right side of the equation above is then
$\int_{0}^{a_{N}}G(x)\,\mathrm{d}x=\int_{0}^{a_{N}}x^{-\alpha}L(x)\,\mathrm{d}x=\frac{1}{1-\alpha}a_{N}^{-\alpha+1}L(a_{N})(1+o(1))=\frac{1}{1-\alpha}\frac{a_{N}}{N}(1+o(1))$
by Karamata’s theorem. (See, e.g., Thm 8.9.2 of [17]). This implies
$N\mathbb{E}[|X_{1}|\mathbbm{1}_{|X_{1}|\leq a_{N}}]/a_{N}\to\frac{1}{1-\alpha}-1.$
Further, from the assumption on $X_{1}$, we can find that
$N\mathbb{E}[X_{1}\mathbbm{1}_{0\leq X_{1}\leq a_{N}}]/a_{N}\to\theta(\frac{1}{1-\alpha}-1).$
We then conclude
$c_{N}/a_{N}\to(2\theta-1)(\frac{1}{1-\alpha}-1),$
which shows that $c_{N}/a_{N}$ converges to a constant depending only on the probability density function. Together with Proposition 3.7, this also shows that $S_{N}/a_{N}$ converges.
@@
See Appendix A for the proof. As a corollary, we can prove the following lemma.
@@
See Appendix A for the proof. The next lemma provides a bound for the number of the eigenvalues whose sizes are comparable to that of the largest eigenvalue.
@@
The eigenvalues $\lambda_{1},..,\lambda_{N}$ of M satisfy
@@
In the following lemma, we show that the limit of $\sum_{i}\lambda_{i}/b_{N}$ is 00.
@@
The eigenvalues $\lambda_{1},..,\lambda_{N}$ of $M$ satisfies